# HACD: Hand-Aware Conditional Diffusion for Monocular Hand-Held Object Reconstruction

### 23 Nov 2023 v1

@@ F-5 $\uparrow$ F-10 $\uparrow$ CD $\downarrow$ F-5 $\uparrow$ F-10 $\uparrow$ CD $\downarrow$
A0 Ours 0.557 0.745 0.265 0.355 0.552 0.659
B0 A0 $\rightarrow$ w/o hand-aware conditioning 0.478 0.667 0.407 0.279 0.461 0.964
B1 A0 $\rightarrow$ w/o hand semantic embedding 0.510 0.694 0.373 0.325 0.496 0.809
B2 A0 $\rightarrow$ w/o hand articulation embedding 0.510 0.693 0.383 0.301 0.483 0.885
B3 A0 $\rightarrow$ w/ GCN hand embedding 0.523 0.713 0.297 0.338 0.537 0.522
B4 A0 $\rightarrow$ w/ global hand embedding 0.524 0.711 0.301 0.309 0.490 0.859
B5 A0 $\rightarrow$ w/o dual-stream conditioning 0.544 0.736 0.274 0.339 0.528 0.756
C0 B0 $\rightarrow$ w/o centroid fixing 0.439 0.609 0.650 0.268 0.452 0.997
C1 B0 $\rightarrow$ w/o centroid prediction 0.317 0.452 2.477 0.229 0.361 1.309
C2 C0 $\rightarrow$ Test with GT object centroid 0.450 0.667 0.362 0.289 0.469 0.932
C3 C0 $\rightarrow$ Test with GT object pose 0.499 0.703 0.343 0.307 0.485 0.841
D0 A0 $\rightarrow$ Predicted hand pose + noise $\sigma=0.1$ 0.556 0.745 0.265 0.348 0.536 0.690
D1 A0 $\rightarrow$ Predicted hand pose + noise $\sigma=0.5$ 0.517 0.707 0.319 0.321 0.504 0.837
E0 A0 $\rightarrow$ GT hand pose 0.596 0.768 0.236 0.372 0.568 0.601
E1 A0 $\rightarrow$ GT hand pose + noise $\sigma=0.1$ 0.572 0.761 0.262 0.358 0.556 0.642
E2 A0 $\rightarrow$ GT hand pose + noise $\sigma=0.5$ 0.535 0.726 0.325 0.314 0.523 0.691
F0 C2 $\rightarrow$ GT object centroid + noise $\sigma=0.1$ 0.442 0.647 0.341 0.281 0.459 1.002
F1 C2 $\rightarrow$ GT object centroid + noise $\sigma=0.5$ 0.411 0.601 0.495 0.236 0.422 1.228
Table 3: Ablation study. (a) Ablation on ObMan [30] and HO3D [29] datasets. (b) Robustness against hand occlusion. We analyze the patterns of the F-5, F-10, and CD metrics as a function of the object visibility ratio on the test set of ObMan [30].
Effectiveness of hand-aware conditioning. In Tab. 3 (a), we show the impact of our proposed hand-aware conditioning. Both the hand semantic embedding (B2 vs. B0, A0 vs. B1) and the geometric hand articulation embedding (B1 vs. B0, A0 vs. B2) result in improved performance. Noteworthy, we can surpass existing methods on the ObMan dataset even without using any hand-aware conditioning (B0).
We further noticed that the way of encoding the hand-aware conditional information has a significant impact on performance. In particular, we have explored two alternative strategies to encode the hand-object interaction. First, inspired by [20], we implement a GCN-based hand embedding, which utilizes a graph convolutional network [23] for feature extraction. We again apply the GCN hand embedding to the partially denoised point cloud at each step. Although the GCN hand embedding implicitly encodes the hand articulation, it does not encode the hand-object interaction. Therefore, simply applying such an embedding as conditioning does not improve performance (See B3 vs. A0). Our hand articulation embedding, on the other hand, applies the unique articulation-aware embedding to each point of the partially denoised point cloud, which explicitly encodes hand-object interaction, thus significantly benefiting the object reconstruction performance. Second, we also utilize standard PointNet [53] to encode 3D hand vertices into a global hand embedding, which is then served to the partially denoised point cloud. Again, without modeling the hand-object interaction, the global hand embedding falls short of providing sufficient semantic information about the hand-occluded object. Consequently, the predictions end up being inferior to our proposed hand semantic embedding (B4 vs. A0).
Effectiveness of dual-stream conditioning. Instead of integrating our unified semantic hand-object and geometric hand articulation embeddings by the proposed dual-stream architecture, we simply concatenate the embeddings. Note that we concatenate a zero matrix to the hand articulation embedding to meet the dimension of unified semantic hand-object embedding. The declined results (B5 vs. A0) demonstrate that dual-stream conditioning enhances reconstruction performance by independently processing semantic and geometric embeddings, ensuring their mutual independence.
Effectiveness of hand-constrained centroid fixing. In Tab. 3 (a), we illustrate the importance of the centroid fixing scheme. Our centroid fixing operation improves the precision and robustness of the semantic and geometric embedding, thus enhancing the performance (B0 vs. C0). Further, without the hand-constrained centroid prediction network, the diffusion model has to simultaneously learn the centroid deviation and object shape, which leads to clearly worse results (C1 vs. B0). Additionally, in C2 we report results when using the actual ground-truth object centroid and in C3 when using the ground-truth object pose. Without our centroid fixing scheme, the results remain non-comparable even when testing with the ground-truth object centroid (C2 vs. B0). This further demonstrates the effectiveness of our proposed centroid fixing scheme.
Robustness against hand pose prediction quality. To demonstrate that our method is able to deal with noisy hand pose estimates, we add various levels of Gaussian noise to predicted hand poses. Notice that adding $\sigma=0.1$ noise has almost no effect on the performance (D0 vs. A0) for both ObMan and HO3D. Further, despite a slight decrease in performance with $\sigma=0.5$ noise (D1 vs. A0), the results clearly show that our approach is robust towards weak hand pose predictions. Additionally, in E0 we explore the upper bound of our method using ground-truth hand poses, and in E1 / E2 we further demonstrate the robustness against noisy hand pose prediction.
Robustness against object centroid prediction quality. We also add different levels of Gaussian noise to the ground-truth object centroid (F0 / F1) to constitute the robustness of our method towards inaccurate object centroid.
@@
Robustness against hand occlusion. To illustrate the robustness of our approach against hand occlusion, we split the test set of ObMan into groups according to the visible ratio of the object and compute the mean F-score and CD for each group. As shown in Tab. 3 (b), when the object is undergoing strong occlusion by the hand (object visible ratio $<$ 50%), iHOI suffers a significant decline in both F-5 and F-10 metrics, while the performance of our approach remains steady. The quantitative results of CD metric also demonstrate that aided by our proposed hand-aware conditioning and hand-constrained centroid fixing scheme, HACD is capable of modeling the uncertainties induced by hand occlusion.
@@ 5 Conclusion
In this paper, we present HACD, a novel hand-aware point cloud diffusion model for single-view hand-held object reconstruction. HACD does not require any object templates, category priors, or depth information, and exhibits superior performance in modeling the uncertainties induced by hand- and self-occlusion. In the core, we introduce hand-aware conditioning to model the semantic and geometric aspects of hand-object interactions. We further propose a hand-constrained centroid-fixing scheme, utilizing the estimated hand vertices to prevent the centroid of the partially denoised point cloud from diverging during diffusion and reverse processes. The experiments on ObMan, HO3D, and MOW datasets demonstrate that our approach surpasses existing methods in both synthetic and real-world scenarios.

### 18 Mar 2024 v2

@@ F-5 $\uparrow$ F-10 $\uparrow$ CD $\downarrow$ F-5 $\uparrow$ F-10 $\uparrow$ CD $\downarrow$
A0 Ours 0.61 0.81 0.11 0.41 0.63 0.34
A1 Ours Oracle 0.67 0.86 0.09 0.51 0.76 0.23
B0 A0 $\rightarrow$ w/o $\mathcal{L}_{mask}$ 0.57 0.76 0.23 0.36 0.56 0.61
C0 B0 $\rightarrow$ w/o dual-stream denoiser 0.54 0.74 0.27 0.34 0.53 0.76
D0 C0 $\rightarrow$ w/o $X_{t}^{HO}$ $X_{t}^{A}$ 0.48 0.67 0.41 0.28 0.46 0.96
D1 C0 $\rightarrow$ w/o $X_{t}^{HO}$ 0.51 0.69 0.37 0.33 0.50 0.81
D2 C0 $\rightarrow$ w/o $X_{t}^{A}$ 0.51 0.69 0.38 0.30 0.48 0.89
D3 C0 $\rightarrow$ w/ GCN hand embedding 0.52 0.71 0.30 0.34 0.53 0.82
D4 C0 $\rightarrow$ w/ global hand embedding 0.52 0.71 0.30 0.31 0.49 0.86
E0 D0 $\rightarrow$ w/o centroid fixing 0.44 0.61 0.65 0.27 0.45 1.00
E1 D0 $\rightarrow$ w/o centroid prediction network 0.32 0.45 2.48 0.23 0.36 1.31
F0 E0 $\rightarrow$ Test with GT object centroid 0.45 0.67 0.36 0.29 0.47 0.93
F1 E0 $\rightarrow$ Test with GT object pose 0.50 0.70 0.34 0.31 0.49 0.84
@@
Robustness against occlusion. To illustrate the robustness of our approach against hand occlusion, we split the test set of ObMan into groups according to the visible ratio of the object and compute the mean F-score for each group. As shown in Tab. 1 (b), when the object is undergoing strong occlusion by the hand (object visible ratio $<$ 50%), iHOI suffers a significant decline in both F-5 and F-10 metrics. Similarly, DDF-HO also experiences an apparent decrease with $<$ 60 % visible ratio. Nevertheless, the performance of our approach instead remains high. Aided by our hand-constrained centroid fixing and the modeling of the hand-induced occlusion, D-SCo exhibits strong robustness against occlusion.
Oracle Experiments. In line with other diffusion models [42, 44] for 3D object reconstruction, in Tab. 4 (A1) we report oracle results for D-SCo. To this end, we predict five possible shapes for each object starting from different sampled Gaussian noises and choose the best sample for each input image with respect to the F-score. Essentially, the oracle results are supposed to demonstrate the probabilistic nature of our method and represent an upper bound. The obtained results underline the ability of D-SCo to overcome the ill-posed essence of the problem. Thanks to the probabilistic formulation in diffusion models, our approach is able to generate multiple plausible shapes (See supplementary material), illustrating our capability of modeling the uncertainty induced by hand- and/or self-occlusion.
The robustness against hand pose prediction quality is further discussed in supplementary material.
@@ 5 Conclusion
In this paper, we present D-SCo, a novel centroid-fixed dual-stream conditional diffusion model for single-view hand-held object reconstruction. D-SCo does not require any object templates, category priors, or depth information, and exhibits superior performance in modeling the uncertainties induced by hand- and self-occlusion. In the core, we propose a hand-constrained centroid-fixing paradigm, utilizing the estimated hand vertices to prevent the centroid of the partially denoised point cloud from diverging during diffusion and reverse processes. Further, a dual-stream denoiser is introduced to semantically and geometrically model hand-object interaction, with a novel unified hand-object semantic embedding enhancing the robustness against occlusion. Our experiments demonstrate that our approach is able to surpass existing methods in both synthetic and real-world scenarios.